Opened 7 years ago

Closed 6 years ago

#462 closed flight processing (fixed)

BGS11/01, flight day 231/2012, Carlsberg

Reported by: knpa Owned by: tipo
Priority: alpha 4 high Milestone: 2012 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by anhi)

Data location: /users/rsg/arsf/arsf_data/2012/flight_data/greenland_iceland/BGS11_01-2012_231_Carlsberg

Data arrived from ARSF via SATA disk on 24/09/2012

BB11/01 Scientific objective:

BB11/01 Priority: a4h

PI: Tapsa Tukiainer

Sensors:

Camera (04/12/2012)
Eagle (27/03/2013)
Hawk (22/02/2013)
Lidar (22/02/2013)

Change History (30)

comment:1 Changed 7 years ago by knpa

  • Description modified (diff)

comment:2 Changed 7 years ago by knpa

  • Description modified (diff)

comment:3 Changed 7 years ago by edfi

Started Navigation

comment:4 Changed 7 years ago by edfi

Navigation Complete, Starting LiDAR Processing

comment:5 Changed 7 years ago by tipo

Started hyperspectral processing.

comment:6 Changed 7 years ago by anhi

Hyperspectral Lines

  • Hawk line 8 (82) crashed at end of line, so is missing the header file, 5gb of data in the raw so worth making a replacement header.
  • Hawk line 12 is recorded between 12:59 and 13:08 so is the cross line (165).
  • Eagle line 12 is recorded between 12:35 and 12:48 so is the line before the cross line (86).
  • Eagle line 13, cross line (165) only has a nav file.
  • Hawk line before cross line is missing (86)

This results in hawk 12 and eagle 12 being different lines (This is all using hdr files and the logsheet).

comment:7 Changed 7 years ago by anhi

RCD Delivery Created

comment:8 Changed 7 years ago by edfi

Delivery Checked RCD some images need removing as they are over exposed. Here are some ranges of images to look at and potentially remove. 01721-01730, 01284-01299, 01935-01946, 01985-01991, 02110-02116, 01656-01668.

comment:9 Changed 7 years ago by tipo

No navigation file found to match times for Hawk line 11. Processing with use nav = false.

comment:10 Changed 7 years ago by tipo

SCT values:

LineEagleHawk
1-0.050.03
2-0.05-0.05
3-0.05-0.05
4-0.05-0.05
5-0.05-0.05
6-0.06-0.05
7-0.05-0.05
8-0.05
9-0.05-0.05
10-0.05-0.05
11-0.052.00
12-0.05-0.05
Last edited 7 years ago by tipo (previous) (diff)

comment:11 Changed 7 years ago by tipo

Hyperspectral processing completed.

comment:12 Changed 7 years ago by anhi

Had to download aster tiles N77W23 to N82W24 to patch onto the LiDAR dem, worth checking the HS dem didn't need these missing tiles when it was made.

comment:13 Changed 7 years ago by anhi

RCD Delivery Created

comment:14 Changed 7 years ago by edfi

Starting Camera delivery check

comment:15 Changed 7 years ago by edfi

Camera Check Complete all ok

comment:16 Changed 7 years ago by anhi

Hyperspectral Delivery check:

  • File identifier mapping table in the Read Me is incorrect, there is a hawk line 165.
  • Two xml files in the project information folder, named differently but look the same.
  • eagle_viewvector_list.txt missing from the sensor_FOV_vectors directory.
  • There is raw data for hawk line 8 which can be used if a hdr file is generated for it.
  • Dropped scans in lines e2, e5, e7, e10, h2, h5, h7, h10, h12.
  • Renamed dem and dem header.
  • Replaced handwritten logsheet with generated logsheet in delivery.

comment:17 Changed 7 years ago by edfi

  • Description modified (diff)

comment:18 Changed 7 years ago by anhi

LiDAR Delivery created:
Pitch: -0.003
Roll: 0.00023

comment:19 Changed 7 years ago by edfi

D/C LiDAR complete. ASCII needed to be converted to DOS friendly which I have done. All else looks fine.

comment:20 Changed 7 years ago by tipo

Made changes to hyperspectral readme.
Spatial binning changes mid project, so two eagle viewvector lists have been created.
Hawk line 8 appears to be nothing but noise.

comment:21 Changed 7 years ago by anhi

Hyperspectral delivery check:

  • aplcorr example command dem name is wrong in the Readme, have corrected in the hyp_genreadme but a new readme needs to be made.

This prevented check_apl_cmd being run so can look at that when it completes but no line changes were made so should be fine.

No other problems

comment:22 Changed 7 years ago by tipo

Corrected readme file.

comment:23 Changed 7 years ago by anhi

Hyperspectral delivery check:

New Readme checked, no problems ready for delivery.

comment:24 Changed 7 years ago by anhi

Delivered LiDAR and Hawk data on 22/02/2013 on disk 151 to Tapani Tukiainen

comment:25 Changed 7 years ago by anhi

  • Description modified (diff)

comment:26 Changed 6 years ago by anhi

Delivered Eagle data on 27/03/2013 on disk 154 to Tapani Tukiainen

comment:27 Changed 6 years ago by anhi

  • Description modified (diff)

comment:28 Changed 6 years ago by tipo

  • Component changed from Processing: general to Archiving
  • Owner set to tipo
  • Status changed from new to assigned

Beginning Archiving

comment:29 Changed 6 years ago by knpa

NEODC currently ingesting.

comment:30 Changed 6 years ago by knpa

  • Resolution set to fixed
  • Status changed from assigned to closed

Backed up by NEODC.

Closing ticket.

Note: See TracTickets for help on using tickets.